# Refined asymptotics for the Cauchy problem for the fast $p$-Laplace evolution equation

### v2

@@ Abstract
Our focus is on the fast diffusion equation driven by the $p$-Laplacian operator, that is $\partial_{t}u=\Delta_{p}u$ with $21<p<$ 2, posed in the whole space ${\mathbb{R}^{N}}$, $N\geq 2$. The nonnegative solutions are expected to converge in time toward a stationary profile. While such convergence had been previously established for $p$ close to $2$, no quantitative rates were known, and the asymptotic behaviour remained poorly understood across the full fast diffusion range. In fact, the long time behaviour of solutions to the $p$-Laplace Cauchy problem drastically change in different subranges of the $p$. Some of them are analysed here for the first time.
In this work, we provide the convergence rates for nonnegative, integrable solutions in the so-called good fast diffusion range, $p_{c}=\tfrac{2N}{N+1}<p<2$, where mass is conserved. We prove that solutions converge to a self-similar profile with matching mass, with explicit rates measured in relative error. Our constructive proof is based on a new entropy method that remains effective even when the entropy is not displacement convex – where optimal transport techniques fail.
In the very fast diffusion range $1<p<p_{c}$, we give the first asymptotic analysis near the extinction time. This regime poses additional challenges: mass is not conserved, solutions vanish in finite time, and no fundamental solutions exist. We uncover new critical exponents – especially in high dimensions – that give rise to markedly different qualitative behaviour depending on the value of $p$.
@@ 1 Introduction
$\begin{cases}\partial_{t}u=\mathrm{div}\left(|\nabla u|^{p-2}\,\nabla u\right)\quad x\in{\mathbb{R}^{N}},\quad t>0\,,\\ u(0,x)={{u_{0}}}(x)\geq 0\,,\quad x\in{\mathbb{R}^{N}}\,,\end{cases}$ (CPLE)
where the exponent $21<p<$ 2. The above equation is called the p-Laplace evolution equation, it is a nonlinear, gradient-driven diffusion equation which is singular in the considered regime $21<p<$ 2, linear when $p=2$, and degenerate for $p>2$. Nonlinear evolution equations involving the $p$-Laplace operator $\Delta_{p}u=\mathrm{div}\left(|\nabla u|^{p-2}\,\nabla u\right)$ have attracted the attention of the mathematical community for over a half of century. The interest come from both the preeminent role of the power-growth operators in the modelling of fluid dynamics [62] and intrinsic mathematical challenges, see e.g. [41, 54, 27, 60] and references therein, and the very recent probabilistic interpretation in terms of a game theoretical approach [34, 66, 64, 65, 63].
@@
The long time behaviour of (CPLE) depends both on the spatial dimension $N$ and the exponent $p$. It is natural, however, to divide analysis into three distinct cases: $21<p<$ 2, $p=2$, and $p>2$. When $p=2$, (CPLE) reduces to the classical heat equation. In the present setting (nonnegative, integrable initial data), the asymptotic behaviour of solutions is well described by the Gaussian profile. We refer to [73, 52] for detailed results on convergence rates and the refined asymptotic analysis in various contexts. When $p>2$, the dynamics differs significantly from the linear case. Notably, solutions exhibit finite speed of propagation: compactly supported initial data generate solutions that remain compactly supported for all times. In this case, the long-time behaviour of solutions is described by the fundamental solution, commonly known as Barenblatt, or Barenblatt-Pattle solution, which is itself compactly supported for all times. Compared to the linear case, the available results are more limited. For a discussion of convergence rates and fine asymptotic behaviour, we refer to [54, 48, 1, 2, 31, 32]. As final comment, we also remark that, in the cases $p=2$ and $p>2$, conservation of mass holds for non-negative $\mathrm{L}^{1}(\mathbb{R}^{N})$ solutions.
The aim of this paper is to present new results and give a broad perspective on the long-time behaviour for (CPLE) in the parameter range $21<p<$ 2. In this regime, as in the linear case, solutions exhibit infinite speed of propagation. However, in contrast with the linear case, they develop fat tails: compactly supported initial data lead to solutions with algebraic spatial decay (power law), rather than a Gaussian profile typical of the linear scenario. The second main distinction from the case $p\geq 2$ is the failure conservation of mass when $p$ is close to $1$. Indeed, for every $t>0$ it holds
@@ 1.1 The good fast diffusion range: $p_{c}<p<2$ and $N\geq 2$
$\mathcal{B}_{M}(t+T,x)=R^{-N}_{T}(t){V_{D}}\left(\frac{x}{R_{T}(t)}\right)\qquad\mbox{and}\qquad D:=\beta^{N\beta\frac{2-p}{p-1}}\,\frac{b_{1}}{M^{\frac{\beta p(2-p)}{p-1}}}\,.$ (1.5)
@@
A natural question arises in view of the above result:
Is it possible to prescribe an explicit rate for the uniform converge in relative $\begin{array}[]{cc}\textit{Is it possible to prescribe an explicit rate for the uniform converge in relative error?}\end{array}$ (Q-1)
@@
This is one of the most delicate points and we need to employ different strategies in different range of parameters since new critical exponents emerge. Thanks to the entropy – entropy production inequality, we deduce (possibly in a weak sense) a differential inequality, that implies the exponential decay of the entropy towards zero:
@@
and the constant $C$ depends on $p$, $N$, ${{v_{0}}}$ and, when $p_{c}<p\leq p_{M}$ also on $D_{1},D_{2}$ of (H). Assume moreover that $0\leq{{v_{0}}}\in C^{2}({\mathbb{R}}^{N})$ is radially symmetric and decreasing ($\partial_{r}{{v_{0}}}\leq 0$), it satisfies (2.1), and that one of the following sets of assumptions holds:
@@
To the best of our knowledge, only four closely related results are present in literature. However, none of them provides a full proof of convergence rates in the entire range $p_{c}<p<2$, whether sharp or not. The first result, valid in the more narrow range $p_{D}<p<2$, is due to Del Pino and Dolbeault [31], and relates the convergence rates to equilibrium to the (possibly optimal) constant in suitable Gagliardo–Nirenberg inequalities, exploiting an entropy method which is analogous to the one we use in the present paper. Unfortunately, in [31] there are no proofs of the announced results. On the other hand, Agueh in [1, 2] provides explicit convergence rates in the same range, and the proof relies on entropy methods based on the Wasserstein gradient flow formulation of the problem. It uses in an essential way displacement convexity of the entropy, so the range of the validity of the method is limited to $p_{D}<p<2$. In the whole range $p_{c}<p<2$, the paper [3] provides convergence rates under similar assumptions. Unfortunately, there are some gaps in the proofs that affect the validity of the results in the range $p_{c}<p<p_{M}$: some integrations by parts are not correct, namely [3, inequality (4.10)] and some quantities that are treated as bounded there, may be unbounded. For instance, the Fisher information happens to be only $\mathrm{L}^{1}$ in time (not necessarily bounded), hence the Gronwall-type argument of [3] fails and it needs to be done differently, e.g. as we do in the proof of Proposition 5.13 using a weak version of Grownwall’s Lemma, see Lemma A.5.
@@
where the infimum is taken among all non-negative and regular enough functions $v$ such that $\int_{\mathbb{R}^{N}}v(y)\,{\rm d}y=\int_{\mathbb{R}^{N}}{V_{D}}(y)\,{\rm d}y$. As we previously explained, in the range $p_{D}\leq p<2$, the value of $\mathcal{C}$ is known and can be deduced from a Gagliardo–Nirenberg–Sobolev-type inequality, see [2, 29]. However, when $p_{c}<p<p_{D}$, this correspondence is not available, and it remains unclear whether $\mathcal{C}\geq 0$ is finite and how to determine its value under general conditions. By taking inspiration from [10, 14], it is possible to compute this constant, at least for radial functions that are close to the Barenblatt profile ${V_{D}}$. Namely, $\Lambda^{\star}$ can be obtained as
@@
Summary. Before proceeding with further results, we believe it is worthwhile to summarise what has been achieved so far in the good fast diffusion range $\frac{2N}{N+1}=p_{c}<p<2$ and clarify under what regime the entropy and the Fisher information are well-defined. Recall that the conservation of mass, i.e. (1.1), is true in this range and the convergence towards the Barenblatt profile ${V_{D}}$ holds as in (1.8), and the parameter $D$ is uniquely determined by the mass of the initial datum ${{v_{0}}}$ as in (1.5). The entropy functional ${\mathcal{E}}[v|{V_{D}}]$, defined in (1.12), is finite along the flow of (R-CPLE) under the sole assumption ${{v_{0}}}\in\mathcal{X}_{p}$ for $p_{M}<p<2$, see Lemma 5.2. It is unbounded in general for $p_{c}<p\leq p_{M}$, but bounded in the range under extra assumption (2.1), see Lemma 5.2. The Fisher information ${\mathcal{I}}[v|{V_{D}}]$, defined in (1.13), is $\mathrm{L}^{\infty}(\tau_{0},\infty)\cap\mathrm{L}^{1}(\tau_{0},\infty)$ for any $\tau_{0}>0$ along the flow of (R-CPLE) in the range $p_{M}<p<2$ under the sole assumption ${{v_{0}}}\in\mathcal{X}_{p}$, see Lemma 5.4. In the same range identity (1.14) holds pointwise. In the range $p_{c}<p\leq p_{M}$ the Fisher information is only in $\mathrm{L}^{1}(\tau_{0},\infty)$ for any $\tau_{0}>0$ along the flow (R-CPLE) and identity (1.14) holds in a weaker sense (5.9), see Lemma 5.5. The relation between the entropy and the Fisher information, i.e. inequality (1.15), is obtained via a Gagliardo–Nirenberg inequality when $p_{D}\leq p<2$, see [2, 29], and by a comparison with a Hardy–Poincaré inequality, see Proposition 5.9 and [28]. We summarise all the above information in the table below.
Range ${\mathcal{E}}[v(\tau)|{V_{D}}]$ ${\mathcal{I}}[v(v)|{V_{D}}]$ $\tfrac{d}{d\tau}{\mathcal{E}}[v(\tau)]=-{\mathcal{I}}[v(v)]$ Inequality (1.15) ${\mathcal{I}}[v(v)|{V_{D}}]\geq c\,{\mathcal{E}}[v(\tau)|{V_{D}}]$ Entropy Decay Rates UCRE Rates
$p_{D}\leq p<2$ $\mathrm{L}^{\infty}(\tau_{0},\infty)$ if ${{v_{0}}}\in\mathcal{X}_{p}$ Lemma 5.2 $\mathrm{L}^{\infty}\cap\mathrm{L}^{1}(\tau_{0},\infty)$ if ${{v_{0}}}\in\mathcal{X}_{p}$ Lemma 5.4 Yes GNS inequality see [2, 29] Theorem 1.17 if ${{v_{0}}}\in\mathcal{X}_{p}$ Optimal in $\mathcal{X}_{p}$: unknown Optimal RD: $\Lambda^{\star}$ Theorem 1.9 if ${{v_{0}}}\in\mathcal{X}_{p}$ Optimal in $\mathcal{X}_{p}$: unknown Optimal expected: $\mathcal{O}(e^{-\tau})$
$p_{M}\leq p<p_{D}$ $\mathrm{L}^{\infty}(\tau_{0},\infty)$ if ${{v_{0}}}\in\mathcal{X}_{p}$ Lemma 5.2 $\mathrm{L}^{\infty}\cap\mathrm{L}^{1}(\tau_{0},\infty)$ if ${{v_{0}}}\in\mathcal{X}_{p}$ Lemma 5.4 Yes Hardy-Poincaré inequality Prop. 5.9 and [28] Theorem 1.17 if ${{v_{0}}}\in\mathcal{X}_{p}$ Optimal in $\mathcal{X}_{p}$: unknown Optimal RD: $\Lambda^{\star}$ Theorem 1.9 if ${{v_{0}}}\in\mathcal{X}_{p}$ Optimal in $\mathcal{X}_{p}$: unknown Optimal expected: $\mathcal{O}(e^{-\frac{\tau}{\beta}})$
$p_{c}<p\leq p_{M}$ $\mathrm{L}^{\infty}(\tau_{0},\infty)$ if (H) Lemma 5.2 $\mathrm{L}^{1}(\tau_{0},\infty)$ if (2.1) Lemma 5.5 As in (5.9) Lemma 5.5 Hardy-Poincaré inequality Prop. 5.9 and [28] Theorem 1.17 if (H) Optimal if (2.1): unknown Optimal RD: $\Lambda^{\star}$ Theorem 1.9 if ${{v_{0}}}\in\mathcal{X}_{p}$ Optimal in $\mathcal{X}_{p}$ or (H): unknown Optimal expected: $\mathcal{O}(e^{-\frac{\tau}{\beta}})$
@@
Is it possible to prescribe an explicit rate for the uniform converge in relative error of radial $\begin{array}[]{c}\textit{Is it possible to prescribe an explicit rate for theuniform converge in relative error of radial derivatives?}\end{array}$ (Q-2)

### v3

@@ Abstract
Our focus is on the fast diffusion equation driven by the $p$-Laplacian operator, that is $\partial_{t}u=\Delta_{p}u$ with $1<p<2$, posed in the whole space $\mathbb{R}^{N}$, $N\geq 2$. The nonnegative solutions, in rescaled variables, are expected to converge in time toward a stationary profile. While such convergence had been previously established for $p$ close to $2$, no quantitative rates were known, and the asymptotic behaviour remained poorly understood across the full fast diffusion range. In fact, the long time behaviour of solutions to the $p$-Laplace Cauchy problem drastically change in different subranges of the $p$. Some of them are analysed here for the first time.
In this work, we provide the convergence rates for nonnegative, integrable solutions in the so-called good fast diffusion range, $p_{c}=\tfrac{2N}{N+1}<p<2$, where mass is conserved. We prove that solutions converge to a self-similar profile with matching mass, with explicit rates measured in relative error. Our constructive proof is based on a new entropy method that remains effective even when the entropy is not displacement convex, where optimal transport techniques fail.
In the very fast diffusion range $1<p<p_{c}$, we give the first asymptotic analysis near the extinction time. This regime poses additional challenges: mass is not conserved, solutions vanish in finite time, and no fundamental solutions exist. We found new critical exponents – especially in high dimensions – that give rise to markedly different qualitative behaviour depending on the value of $p$.
@@ 1 Introduction
[…]
where the exponent $1<p<2$. The above equation is called the p-Laplace evolution equation, it is a nonlinear, gradient-driven diffusion equation which is singular in the considered regime $1<p<2$, linear when $p=2$, and degenerate for $p>2$. Nonlinear evolution equations involving the $p$-Laplace operator $\Delta_{p}u=\mathrm{div}\left(|\nabla u|^{p-2}\,\nabla u\right)$ have attracted the attention of the mathematical community for over a half of century. The interest come from both the preeminent role of the power-growth operators in the modelling of fluid dynamics [63] and intrinsic mathematical challenges, see e.g. [42, 55, 27, 61] and references therein, and the very recent probabilistic interpretation in terms of a game theoretical approach [35, 67, 65, 66, 64].
@@
The long time behaviour of (CPLE) depends both on the spatial dimension $N$ and the exponent $p$. It is natural, however, to divide analysis into three distinct cases: $1<p<2$, $p=2$, and $p>2$. When $p=2$, (CPLE) reduces to the classical heat equation. In the present setting (nonnegative, integrable initial data), the asymptotic behaviour of solutions is well described by the Gaussian profile. We refer to [74, 53] for detailed results on convergence rates and the refined asymptotic analysis in various contexts. When $p>2$, the dynamics differs significantly from the linear case. Notably, solutions exhibit finite speed of propagation: compactly supported initial data generate solutions that remain compactly supported for all times. In this case, the long-time behaviour of solutions is described by the fundamental solution, commonly known as Barenblatt, or Barenblatt-Pattle solution, which is itself compactly supported for all times. Compared to the linear case, the available results are more limited. For a discussion of convergence rates and fine asymptotic behaviour, we refer to [55, 49, 1, 2, 31, 32]. As final comment, we also remark that, in the cases $p=2$ and $p>2$, conservation of mass holds for non-negative $\mathrm{L}^{1}(\mathbb{R}^{N})$ solutions.
The aim of this paper is to present new results and give a broad perspective on the long-time behaviour for (CPLE) in the parameter range $1<p<2$. In this regime, as in the linear case, solutions exhibit infinite speed of propagation. However, in contrast with the linear case, they develop fat tails: compactly supported initial data lead to solutions with algebraic spatial decay (power law), rather than a Gaussian profile typical of the linear scenario. The second main distinction from the case $p\geq 2$ is the failure conservation of mass when $p$ is close to $1$. Indeed, for every $t>0$ it holds
@@ 1.1 The good fast diffusion range: $p_{c}<p<2$ and $N\geq 2$
$\mathcal{B}_{M}(t+T,x)=R^{-N}_{T}(t){V_{D}}\left(\frac{x}{R_{T}(t)}\right)\qquad\mbox{and}\qquad D:=\beta^{N\beta\frac{2-p}{p-1}}\,\frac{b_{1}}{M^{p\beta\frac{2-p}{p-1}}}\,.$ (1.5)
@@
A natural question arises in the view of the above result:
[…]
@@
This represents one of the most delicate aspects of the problem, since the emergence of new critical exponents necessitates different strategies across distinct parameter ranges. Thanks to the entropy – entropy production inequality, we deduce (possibly in a weak sense) a differential inequality, that implies the exponential decay of the entropy towards zero:
@@
and the constant $C$ depends on $p$, NN, $v_{0}$}} and, when $p_{c}<p\leq p_{M}$ also on $D_{1},D_{2}$ of (H). Assume moreover that $0\leq{{v_{0}}}\in C^{2}({\mathbb{R}}^{N})$ is radially symmetric and decreasing ($\partial_{r}{{v_{0}}}\leq 0$), it satisfies (2.1), and that one of the following sets of assumptions holds:
@@
To the best of our knowledge, only a few closely related results are present in literature, and none of them provides a proof of convergence rates in the entire range $p_{c}<p<2$, whether sharp or not. The first pioneering result, valid in the range $p_{D}<p<2$, is due to Del Pino and Dolbeault, announced in [31] and proven in detail in [32]. Their method connects convergence rates to equilibrium with the optimal constant in suitable Gagliardo–Nirenberg inequalities, obtained in [33], using a new entropy method, that inspired the one developed in the present paper. On the other hand, Agueh in [1, 2] provides explicit convergence rates in the same range, and the proof relies on entropy methods based on the Wasserstein gradient flow formulation of the problem. It uses in an essential way displacement convexity of the entropy, so the range of the validity of the method is limited to $p_{D}<p<2$. In the whole range $p_{c}<p<2$, the paper [3] provides convergence rates under similar assumptions. Unfortunately, there are some gaps in the proofs that affect the validity of the results in the range $p_{c}<p<p_{M}$: some integrations by parts are not correct, namely [3, inequality (4.10)] and some quantities that are treated as bounded there, may be unbounded. For instance, the Fisher information happens to be only $\mathrm{L}^{1}$ in time (not necessarily bounded), hence the Gronwall-type argument of [3] fails and it needs to be done differently, e.g. as we do in the proof of Proposition 5.13 using a weak version of Grownwall’s Lemma, see Lemma A.5.
@@
where the infimum is taken among all non-negative and regular enough functions $v$ such that $\int_{\mathbb{R}^{N}}v(y)\,{\rm d}y=\int_{\mathbb{R}^{N}}{V_{D}}(y)\,{\rm d}y$. As we previously explained, the value of $\mathcal{C}$ in the range $p_{D}\leq p<2$ is known, as it is the optimal constant in a suitable Gagliardo–Nirenberg–Sobolev-type inequality, whose value was obtained for the first time in [33], see also [29] for an alternative proof. However, when $p_{c}<p<p_{D}$, this correspondence is not available, and it remains unclear whether $\mathcal{C}\geq 0$ is finite and how to determine its value under general conditions. By taking inspiration from [10, 14], it is possible to compute this constant, at least for radial functions that are close to the Barenblatt profile $V_{D}$}. Namely, $\Lambda^{\star}$ can be obtained as
@@
Before proceeding with further results, we believe it is worthwhile to summarise what has been achieved so far in the good fast diffusion range $\frac{2N}{N+1}=p_{c}<p<2$ and clarify under what regime the entropy and the Fisher information are well-defined. Recall that the conservation of mass, i.e. (1.1), is true in this range and the convergence towards the Barenblatt profile $V_{D}$} holds as in (1.8), and the parameter $D$ is uniquely determined by the mass of the initial datum $v_{0}$}} as in (1.5). The entropy functional ${\mathcal{E}}[v|{V_{D}}]$, defined in (1.12), is finite along the flow of (R-CPLE) under the sole assumption ${{v_{0}}}\in\mathcal{X}_{p}$ for $p_{M}<p<2$, see Lemma 5.2. It is unbounded in general for $p_{c}<p\leq p_{M}$, but bounded in the range under extra assumption (2.1), see Lemma 5.2. The Fisher information ${\mathcal{I}}[v|{V_{D}}]$, defined in (1.13), is $\mathrm{L}^{\infty}(\tau_{0},\infty)\cap\mathrm{L}^{1}(\tau_{0},\infty)$ for any $\tau_{0}>0$ along the flow of (R-CPLE) in the range $p_{M}<p<2$ under the sole assumption ${{v_{0}}}\in\mathcal{X}_{p}$, see Lemma 5.4. In the same range identity (1.14) holds pointwise. In the range $p_{c}<p\leq p_{M}$ the Fisher information is only in $\mathrm{L}^{1}(\tau_{0},\infty)$ for any $\tau_{0}>0$ along the flow (R-CPLE) and identity (1.14) holds in a weaker sense (5.9), see Lemma 5.5. The relation between the entropy and the Fisher information, i.e. inequality (1.15), is obtained via a Gagliardo–Nirenberg inequality when $p_{D}\leq p<2$, see [2, 29], and by a comparison with a Hardy–Poincaré inequality, see Proposition 5.9 and [28]. We summarise all the above information in the table below.
Range ${\mathcal{E}}[v(\tau)|{V_{D}}]$ ${\mathcal{I}}[v(v)|{V_{D}}]$ $\tfrac{d}{d\tau}{\mathcal{E}}[v(\tau)]=-{\mathcal{I}}[v(v)]$
Inequality (1.15)
${\mathcal{I}}[v(v)|{V_{D}}]\geq c\,{\mathcal{E}}[v(\tau)|{V_{D}}]$
Entropy Decay
Rates
UCRE
Rates
$p_{D}\leq p<2$
$\mathrm{L}^{\infty}(\tau_{0},\infty)$
if ${{v_{0}}}\in\mathcal{X}_{p}$
Lemma 5.2
$\mathrm{L}^{\infty}\cap\mathrm{L}^{1}(\tau_{0},\infty)$
if ${{v_{0}}}\in\mathcal{X}_{p}$
Lemma 5.4
Yes
GNS inequality
see [2, 29]
Theorem 1.17 if ${{v_{0}}}\in\mathcal{X}_{p}$
Optimal in $\mathcal{X}_{p}$: unknown
Optimal RD: $\Lambda^{\star}$
Theorem 1.9 if ${{v_{0}}}\in\mathcal{X}_{p}$
Optimal in $\mathcal{X}_{p}$: unknown
Optimal expected: $\mathcal{O}(e^{-\tau})$
$p_{M}\leq p<p_{D}$
$\mathrm{L}^{\infty}(\tau_{0},\infty)$
if ${{v_{0}}}\in\mathcal{X}_{p}$
Lemma 5.2
$\mathrm{L}^{\infty}\cap\mathrm{L}^{1}(\tau_{0},\infty)$
if ${{v_{0}}}\in\mathcal{X}_{p}$
Lemma 5.4
Yes
Hardy-Poincaré inequality
Prop. 5.9 and [28]
Theorem 1.17 if ${{v_{0}}}\in\mathcal{X}_{p}$
Optimal in $\mathcal{X}_{p}$: unknown
Optimal RD: $\Lambda^{\star}$
Theorem 1.9 if ${{v_{0}}}\in\mathcal{X}_{p}$
Optimal in $\mathcal{X}_{p}$: unknown
Optimal expected: $\mathcal{O}(e^{-\frac{\tau}{\beta}})$
$p_{c}<p\leq p_{M}$
$\mathrm{L}^{\infty}(\tau_{0},\infty)$
if (H)
Lemma 5.2
$\mathrm{L}^{1}(\tau_{0},\infty)$
if (2.1)
Lemma 5.5
As in (5.9)
Lemma 5.5
Hardy-Poincaré inequality
Prop. 5.9 and [28]
Theorem 1.17 if (H)
Optimal if (2.1): unknown
Optimal RD: $\Lambda^{\star}$
Theorem 1.9 if ${{v_{0}}}\in\mathcal{X}_{p}$
Optimal in $\mathcal{X}_{p}$ or (H): unknown
Optimal expected: $\mathcal{O}(e^{-\frac{\tau}{\beta}})$
@@
[…]